OpenStudy (anonymous):

Topic: Base e and Natural Logarithms ln (x-2) = 2 and another one is ln x + ln 2x = 2 @whpalmer4

OpenStudy (whpalmer4):

Okay, are you supposed to solve for \(x\)?

OpenStudy (whpalmer4):

\[\ln x = a\]means that \[e^a = x\] So if we want to solve my equation for \(x\), we raise \(e\) to the power of both sides: \[e^{\ln x} = e^a\]but \[b^{\log_b x} = b\] so the left side simplifies: \[e^{\ln x} = e^a\]\[x = e^a\]

OpenStudy (whpalmer4):

With that knowledge, you should be able to solve the first problem. Do so and I'll check your answer.

OpenStudy (anonymous):

so x-2=e^2

OpenStudy (whpalmer4):

Yes, although I would write \[x = 2+e^2\] if you are asked to solve for \(x\)

OpenStudy (anonymous):

oh ok I forgot about the 2

OpenStudy (whpalmer4):

Now for the second one, you're going to use a handy property of logarithms: \[\log a + \log b = \log (a*b)\]

OpenStudy (anonymous):

wait what button do I press on my calculator for e

OpenStudy (whpalmer4):

I would just leave the answer as you have it unless specifically asked for a decimal number. You may have an \(e^x\) button on your calculator which calculates \(e^x\) when you've entered \(x\) already... what model calculator?

OpenStudy (whpalmer4):

\[e\approx 2.7182818284590452354\]

OpenStudy (anonymous):

TI-84 plus

OpenStudy (anonymous):

but I got 2.6931

OpenStudy (whpalmer4):

looks like it might be the second function on the \(\ln\) button

OpenStudy (anonymous):

oh ok I got it I have to press 2nd then the button

OpenStudy (whpalmer4):

Yeah, 0.6931 is the natural log of 2. One of the values I've memorized :-)

OpenStudy (anonymous):

so 9.3890 is the answer

OpenStudy (whpalmer4):

it turns out that you can memorize just a handful of values and estimate many from that. for example, \(\ln 8 = \ln 2^3 = 3\ln 2 = 3*0.693 \approx 2.1\) (true value is 2.07944...) Yes, 9.38906... is your value of \(x\)

OpenStudy (anonymous):

I see there are multiple ways they can be put :)

OpenStudy (whpalmer4):

but as I suggested, as long as you know how to convert it to a number, it's generally more useful to keep it in the "symbolic" (and exact) form \(x = 2+e^2\)

OpenStudy (whpalmer4):

You can always turn it into a number, but recognizing that 9.38906... is \(2 + e^2\) is much more challenging :-)

OpenStudy (anonymous):

it is but if I do get those numbers down it would have to be when I'm not busy lol

OpenStudy (anonymous):

would the second problem be ln 2x^2

OpenStudy (whpalmer4):

My point is that in general, keeping things in forms like \(35\pi, e^2\), etc. allows you to have more information about what you are dealing with, and possibly recognize paths to a solution which you wouldn't if you just had a number. Yes, the second problem becomes \[\ln 2x^2 = 2\]Can you solve that for \(x\)?

OpenStudy (anonymous):

I'm not really sure what to do :(

OpenStudy (whpalmer4):

Well, again, raise e to both sides: \[e^{\ln 2x^2} = e^2\]Now simplify that the same way you did with the first one

OpenStudy (anonymous):

wait again where did you get the e from and why di you put it in. Was it to cancel out the ln?

OpenStudy (whpalmer4):

e is the base of the natural logarithm

OpenStudy (whpalmer4):

\[\ln x = \log_e x\]

OpenStudy (whpalmer4):

it's such a useful concept that they invented that shorthand...

OpenStudy (whpalmer4):

you haven't had enough math yet to see it, but it turns out that once you get into calculus and beyond, \(e\) pops up everywhere you look...there's even a book called "E: the story of a number"

OpenStudy (anonymous):

lol really that's fuuny

OpenStudy (anonymous):

now I'm at 2x^2+ 7.38905....

OpenStudy (whpalmer4):

There's an elegant identity that links \(e\) and \(\pi\) and \(i = \sqrt{-1}\) and \(1\) and \(0\) together: \[e^{i\pi} +1 = 0\]

OpenStudy (anonymous):

what is that formula called?

OpenStudy (whpalmer4):

Please, you shouldn't have any decimal numbers in sight while working these problems... \[e^{\ln 2x^2} = e^2\]\[2x^2 = e^2\]Now continue...

OpenStudy (whpalmer4):

https://en.wikipedia.org/wiki/Euler's_identity

OpenStudy (anonymous):

this might sound dumb but what do I do with the 2x^2

OpenStudy (whpalmer4):

Okay, you're trying to get the left side to look like \{x = \]right?

OpenStudy (anonymous):

ummhmmm

OpenStudy (whpalmer4):

sorry, \[x=\]

OpenStudy (whpalmer4):

but you have \[2x^2 = \]

OpenStudy (anonymous):

do I divide

OpenStudy (whpalmer4):

why don't you start by dividing both sides by 2?

OpenStudy (anonymous):

oh ok lol

OpenStudy (whpalmer4):

then take the square root of both sides

OpenStudy (whpalmer4):

"brain to forehead: look out forehead, here comes palm!" :-)

OpenStudy (anonymous):

lol never heard that one before, so now I;kmkn

OpenStudy (anonymous):

sorry something is going on with my backspace button

OpenStudy (whpalmer4):

another good book: "Dr. Euler's Fabulous Formula: Cures Many Mathematical Ills" :-)

OpenStudy (whpalmer4):

yeah, I sometimes get OpenStudy in a state where the backspace key doesn't work. annoying!

OpenStudy (anonymous):

I well any how how would I squareroot 2e^2

OpenStudy (anonymous):

Your still there right

OpenStudy (whpalmer4):

Yes. \[2x^2 =e^2\]Divide both sides by 2 \[\frac{2x^2}{2} = \frac{e^2}{2}\]\[x^2=\frac{e^2}{2}\] Take square root of both sides

OpenStudy (anonymous):

????? I'm dead

OpenStudy (whpalmer4):

Why? What is the square root of \(x^2\)?

OpenStudy (anonymous):

x lol I got that one

OpenStudy (anonymous):

oh so its 2

OpenStudy (whpalmer4):

what is 2?

OpenStudy (anonymous):

the answer right

OpenStudy (whpalmer4):

I knew when I asked that you were going to say that :-(

OpenStudy (anonymous):

really lol y?

OpenStudy (whpalmer4):

but I was hoping that you'd tell me something else, like "that's the number of problems I have left" because it is not the answer to this problem

OpenStudy (anonymous):

lol omg I'm never going to get this

OpenStudy (whpalmer4):

\[x^2 = \frac{e^2}{2}\]\[\sqrt{x^2} = \sqrt{\frac{e^2}{2}} = \frac{\sqrt{e^2}}{\sqrt{2}}\]right?

OpenStudy (whpalmer4):

then simplifying (assuming that \(x>0\)): \[x = \frac{e}{\sqrt{2}}\]

OpenStudy (whpalmer4):

actually, no need to assume, we can see that \(x > 0\) here. but in general, you can't just simplify \(\sqrt{x^2}\rightarrow x\)unless you know that \(x>0\)

OpenStudy (anonymous):

ok I get it bu tdo you mind helping me with another problem plzzzzz

OpenStudy (whpalmer4):

How could I possibly refuse? :-)

OpenStudy (anonymous):

OMG thanx ok here we goo>>> ln (x+3) = 4

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

:-( I'm happy to help, but I do ask that you try to recognize problems of a sort we've already worked out...

OpenStudy (anonymous):

ok lol so you want me to do this one on my own? :)

OpenStudy (whpalmer4):

I think you can, yes. Happy to check the result for you...

OpenStudy (whpalmer4):

Ever heard of the surgeon's motto? "See one, do one, teach one" :-)

OpenStudy (anonymous):

ok hopefully I can DO ONE lol

OpenStudy (anonymous):

kk first you take e and square it to ln in which they cancel. Then you take the e and square it to 4

OpenStudy (anonymous):

51.5982 would be the answer rounded am I correct

OpenStudy (whpalmer4):

could we have that in plain mathese, please? :-) That's the correct (approximate) number. But really, I'd prefer \[x = e^4-3\]which is exact.

OpenStudy (anonymous):

ok then lol :p

OpenStudy (anonymous):

you think you could help me with even more problems

OpenStudy (whpalmer4):

\(e\) is irrational, so it is not possible to write down the exact string of digits which represents it, because it goes on forever...

OpenStudy (anonymous):

but our teacher wants us too so Idk

OpenStudy (whpalmer4):

\(\pi\) is also irrational (meaning it can't be represented as the ratio of two integers) Oh, if your teacher wants a decimal, then by all means, that's what you should provide. I thought I said "unless the problem asks for a decimal" when I first expressed my preference for the symbolic, exact form.... I can't be held responsible if your teacher prefers inexact answers that take up more space :-)

OpenStudy (anonymous):

lol I guess and r u ready for properties of logs or do I HAVE YOU TIRE D YET

OpenStudy (anonymous):

TIRED

OpenStudy (whpalmer4):

let's do another, if you're up for it...

OpenStudy (whpalmer4):

but why don't you post it in a fresh question...

OpenStudy (anonymous):

@whpalmer4 Use log\[\log_{2} 3 \approx 1.5850 and \log_{2} 5\approx 2.3219 \to approximate the value of each expression.\]

OpenStudy (whpalmer4):

two reasons: 1) you can only give a single medal per question, and 2) when I flip back to this page, I have to scroll to the bottom each time, which is getting successively more tedious :-)

OpenStudy (anonymous):

1) \[\log_{2} 25\]

OpenStudy (anonymous):

oh ok hold then lol